OpenStudy (anonymous):

What is the equation of the line that passes through (–2, 3) and is parallel to 2x + 3y = 6?

OpenStudy (anonymous):

let equation of line is y=mx+c you know slope m by given equation of line 2x + 3y = 6 m=-2/3 by putting value of m you will get y=(-2/3)x+c now put point (-2,3) in this line and get c and after put value of c in same equation

OpenStudy (anonymous):

sooo whats the answer

OpenStudy (anonymous):

please calculate it by own and you tell me final answer.

OpenStudy (anonymous):

im trying to get this answer quick i have a time limit please tell me

OpenStudy (anonymous):

continuing from above y=(-2/3)x+c now put point (-2,3) in this line and get c so value of c =5/3 now equan is 2x+3y=5